### v1

@@ 1 Introduction
The $T\bar{T}$ deformation [1, 2, 3] provides an instructive case of an irrelevant deformation of quantum field theory (QFT) that remains well-defined and exactly solvable. One of the most illuminating perspectives on this deformation is its interpretation as a coupling to two-dimensional gravity [4, 5, 6]. Within this framework, $T\bar{T}$-deformed QFTs can be reformulated as quantum field theories on fluctuating geometries governed by topological or massive gravity in two dimensions.
@@ 3 A constructive approach to the gravitational dual – $T{\bar{T}}$ braneworld holography
where central charge $c$ is related to the dimensionless Newton constant $G$ through the standard holographic dictionary, $c=3/(2G)$, and the trace condition has been used. [^5] In connection with Footnote 4, the presence of matter sources on the boundary leads to additional contributions to the conformal anomaly [26].
@@ 4.4 All-order proof
$\qquad=\left(\delta_{i}^{k}-2\mu T_{\rm cft}(\delta)_{i}^{k}+\mu^{2}T^{2}_{\rm cft}(\delta)_{k}^{i}\right)\left(1-\mu T_{\rm cft}(\delta)\right)^{-2}_{kj}\,+\,{\rm derivative\,\,terms}$
@@
which reproduces (4.6), (4.12), and (4.22). [^7] We used $\delta^{a}_{i}\mathop{\mathrm{Tr}}\nolimits T_{\rm cft}^{2}(\delta)=2\delta^{a}_{k}\mathop{\mathrm{Tr}}\nolimits T_{\rm cft}^{2}(\delta)^{k}_{i}$, $2T_{\rm cft}^{3}(\delta)^{k}_{i}=T_{\rm cft}^{2}(\delta)T_{\rm cft}(\delta)^{k}_{a}$, and $\mathop{\mathrm{Tr}}\nolimits T_{\rm cft}^{3}(\delta)=0$.
@@
As a byproduct of our analysis, combining (2.5) and (2.8), we obtain an all-order expression for the $T\bar{T}$-deformed stress tensor on $\mathbb{R}^{2}$ in the saddle-point approximation:
@@
where $X^{i}_{j}=T_{\rm cft}(\delta)^{i}_{j}+{1\over 16\pi G}R(e^{\ast})^{i}_{j}$. The stress tensor with spacetime indices takes the similar form:
@@
where $f^{a}_{j}=\delta^{a}_{j}$ on $\mathbb{R}^{2}$. One caveat is that, since the curvature $R$ is a functional of $(e^{\ast})^{a}_{i}$, obtaining a closed expression purely in terms of the flat-space CFT stress tensor $T_{\rm cft}(\delta)$ still requires solving either the saddle-point equation (4.35) or the all-order zweibein expression (4.34) for $(e^{\ast})^{a}_{i}$.

### v2

@@ 1 Introduction
The $T\bar{T}$ deformation [1, 2, 3] provides an instructive case of an irrelevant deformation of quantum field theory (QFT) that remains well-defined and exactly solvable. [^1] A recent comprehensive review can be found in [4]. One of the most illuminating perspectives on this deformation is its interpretation as a coupling to two-dimensional gravity [5, 6, 7]. Within this framework, $T\bar{T}$-deformed QFTs can be reformulated as quantum field theories on fluctuating geometries governed by topological or massive gravity in two dimensions.
@@ 3 A constructive approach to the gravitational dual – $T{\bar{T}}$ braneworld holography
where central charge $c$ is related to the dimensionless Newton constant $G$ through the standard holographic dictionary, $c=3/(2G)$, and the trace condition has been used. [^6] In connection with Footnote 5, the presence of matter sources on the boundary leads to additional contributions to the conformal anomaly [27].
@@ 4.4 All-order proof
$\qquad=\left(\delta_{i}^{k}-2\mu T_{\rm cft}(\delta)_{i}^{k}+\mu^{2}T^{2}_{\rm cft}(\delta)^{k}_{i}\right)\left(1-\mu T_{\rm cft}(\delta)\right)^{-2}_{kj}\,+\,{\rm derivative\,\,terms}$
@@
which reproduces (4.6), (4.12), and (4.22). [^8] We used $\delta^{a}_{i}\mathop{\mathrm{Tr}}\nolimits T_{\rm cft}^{2}(\delta)=2\delta^{a}_{k}\mathop{\mathrm{Tr}}\nolimits T_{\rm cft}^{2}(\delta)^{k}_{i}$, $2T_{\rm cft}^{3}(\delta)^{k}_{i}=\mathop{\mathrm{Tr}}\nolimits T_{\rm cft}^{2}(\delta)T_{\rm cft}(\delta)^{k}_{a}$, and $\mathop{\mathrm{Tr}}\nolimits T_{\rm cft}^{3}(\delta)=0$.
@@
5 Derivation of the $T\bar{T}$ deformation as dynamical coordinate transformations
To further strengthen the foundation of our braneworld approach to $T\bar{T}$ holography, we present a derivation that interprets the $T\bar{T}$ deformation as arising from dynamical coordinate transformations [18, 19]. (See also related discussions in [11].)
First, observe that the induced metric on a generic constant radial slice takes the particularly simple form
$ds_{\rm ind}^{2}=g^{\rm ind}_{ij}(e^{\ast})dx^{i}dx^{j}=\left((1-4\pi G\rho^{2}X)^{2}\right)_{i}^{k}\left(1-\mu X\right)^{-2}_{kj}dx^{i}dx^{j}\ .$ (5.1)
This structure suggests that the effect of the $T\bar{T}$ deformation can be reinterpreted as a coordinate transformation, $Y^{i}\mapsto x^{i}$:
$dY^{i}=\left((1-\mu X)^{-1}\right)^{i}_{j}dx^{j}\ .$ (5.2)
Under this map, and using the commuting property of the symmetric matrix $M(\alpha)\equiv 1-\alpha X$, namely $M(\alpha)M(\beta)=M(\beta)M(\alpha)$, the induced metric takes the initial Fefferman-Graham form:
$ds_{\rm ind}^{2}=(1-4\pi G\rho^{2}X)^{2}_{ij}dY^{i}dY^{j}\ ,$ (5.3)
demonstrating that the $T\bar{T}$ deformation is fully absorbed into the redefined coordinates $Y^{i}$.
We now establish that the coordinate transformation (5.2) is precisely the dynamical coordinate transformation proposed in [18], which generalizes the infinitesimal version of [19]: As a byproduct of our analysis in the previous section, combining (2.5) and (2.8), we obtain an all-order expression for the $T\bar{T}$-deformed stress tensor on $\mathbb{R}^{2}$ in the saddle-point approximation:
@@ 5 Derivation of the $T\bar{T}$ deformation as dynamical coordinate transformations
where $X^{i}_{j}=T_{\rm cft}(\delta)^{i}_{j}+{1\over 16\pi G}R(e^{\ast})^{i}_{j}$. The corresponding form with spacetime indices is
@@
where $f^{a}_{j}=\delta^{a}_{j}$ on $\mathbb{R}^{2}$. A caveat is that since the curvature $R$ is a functional of $(e^{\ast})^{a}_{i}$, obtaining a closed expression purely in terms of the flat-space CFT stress tensor $T_{\rm cft}(\delta)$ still requires solving either the saddle-point equation (4.35) or the all-order zweibein relation (4.34). From this expression one can establish the identity
$\left(1-\mu T(e^{\ast})\right)^{i}_{j}={(1-\mu X)^{i}_{j}\over\det(1-\mu X)}\ ,$ (5.6)
using the fact that $\left((1-\mu X)^{-1}\right)^{i}_{j}=\frac{\delta^{i}_{k}+\mu(X^{i}_{k}-\delta^{i}_{k}\mathop{\mathrm{Tr}}\nolimits X)}{\det(1-\mu X)}$ for a symmetric matrix $X$. This further implies that
$\epsilon^{ik}\epsilon_{jl}\left(1-\mu T(e^{\ast})\right)^{l}_{k}={\epsilon^{ik}\epsilon_{jl}(1-\mu X)^{l}_{k}\over\det(1-\mu X)}=\left((1-\mu X)^{-1}\right)^{i}_{j}\ ,$ (5.7)
since the contraction of two Levi-Civita symbols exchanges indices $1$ and $2$ and flips the sign of off-diagonal components.
Thus, the coordinate transformation (5.2) can be written equivalently as
$dY^{i}=dx^{i}-\mu\epsilon^{ik}\epsilon_{jl}T(e^{\ast})^{l}_{k}dx^{j}\ .$ (5.8)
This completes the holographic derivation of the dynamical coordinate transformation, extending the infinitesimal version of [19, 21] to its finite form [18].
As a final remark, we note that, using the identity (5.7), the deformation of the boundary metric can be expressed as
$g^{(0)}(e^{\ast})^{i}_{j}=\delta^{i}_{j}-2\mu\left(\delta^{i}_{j}\,\mathop{\mathrm{Tr}}\nolimits\,T(e^{\ast})-T(e^{\ast})^{i}_{j}\right)+\mu^{2}\left(\delta^{i}_{j}\,\mathop{\mathrm{Tr}}\nolimits\,T(e^{\ast})^{2}-T^{2}(e^{\ast})^{i}_{j}\right)\ .$ (5.9)
This structure resembles the “mixed boundary conditions” of [28], but with an essential difference: here the stress tensor is the $T\bar{T}$-deformed one, rather than that of the undeformed CFT.